Circular No. 8264

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(121) HERMIONE F. Marchis and C. Laver, University of California at Berkeley (UCB);

and J. Berthier, Institut de Mécanique Céleste et de Calcul des Ephémerides (IMCCE) — in collaboration with P. Descamps and D. Hestroffer, IMCCE; and I. de Pater, UCB — report evidence of a bilobated shape of this mainbelt minor planet from observations performed on 2003 Dec. 6.521 and Dec. 7.528 UT at 2.2 μm using the Keck 10-m telescope (+ NIRC-2) and its adaptive-optics system. The ‘peanut’-shaped minor planet appears 230 km (0′′.126) long with two lobes of radius 60 and 50 km separated by 120 km and perhaps linked by a bridge of matter 80 km wide, or as two connected components of radius 90 and 60 km (a ‘snowman’ shape) separated by a center-to-center distance of 115 km. Angular resolution provided on these images cannot discern between these models but the images clearly rule out a simple ellipsoid shape. Additionally, the comparison with the physical ephemeris indicates a pole solution that agrees with de Angelis (1995, Planet. Space Sci. 43, 649) and confirms the rotation rate of 5.551 hr measured recently by R. Behrend (Geneva Observatory) on the basis of new photometric observations made by R. Roy (Blauvac, France) and P. Baudoin (Le Havre, France). Note that (121) is also known to have a moonlet companion orbiting at 790 km around the primary (cf. IAUC 7980). An occultation involving this extremely interesting target will occur on 2004 Feb. 16 UT and will be visible mainly from Europe. Adaptive-optics images and information about the occultation are available at the website

http://astron.berkeley.edu/~fmarchis/Science/HermioneKeck/.

COMETS C/2002 X7, X8, X9, X10, X11, X12 (SOHO) Further to IAUC 8263, K. Battams reports measures of additional

Kreutz comets found by X.-m. Zhou (C/2002 X7, X10, and X12), M. Meyer (C/2002 X8), S. Hönig (C/2002 X9), and R. Kracht (C/2002 X11) on SOHO C2 images (astrometry and orbital elements on MPEC 2004-A36).

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 X7 Dec. 1.329  16^{h}30.^{m}3  −23^{o}30′
C/2002 X8  5.271  16 48.3  −24 09
C/2002 X9  5.288  16 50.7  −24 18
C/2002 X10  5.564  16 49.7  −24 10
C/2002 X11  8.963  17 05.7  −24 37
C/2002 X12  9.564  17 08.2  −24 38
2004 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green